Regular marbles sell at 100 for $2.00. Large marbles (shooters) sell at 25 for $2.00. To the nearest whole marble, how many of each kind should be used to make a package of 100 marbles selling for $3?
Answer by stanbon(75887)

Regular marbles sell at 100 for $2.00. Large marbles (shooters) sell at 25 for $2.00. To the nearest whole marble, how many of each kind should be used to make a package of 100 marbles selling for $3?
---------------------------
Regular marbles cost $0.02 a piece.
Shooters cost $0.08 a piece.
--------------------------
Let # of regular marbles be "x" ; value of these is 0.02x dollars
---------------------------
# of shooters is "100-x" ; value of these is 0.08(100-x)= 8-0.08x dollars
----------------------------
# in mixture is 100 ; value = $3
----------------------------------
EQUATION:
value + value = value
0.02x + 8-0.08x = 3
-0.06x = -5
x = 83 1/3 (number of regular marbles)
100-x = 16 2/3 ( number of shooters)
----------------------------------------
Comment: The value of x and 100-x should turn out to be whole numbers.
Check your problem statement to see if you have the correct data.
==========================================
Cheers,
Stan H.